OpenStudy (anonymous):

Choose the equation of the line passing through the point (4, -2) and perpendicular to y = x - 7. y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

hero (hero):

What about the previous question?

OpenStudy (anonymous):

I already did it

OpenStudy (anonymous):

but i really need this one

hero (hero):

Using y = mx + b, input x = 4, y = -2, and m = -1 into the formula to find b: -2 = (-1)(4) + b -2 = -4 + b 4 - 2 = b 2 = b Since m = -1 and b = 2, the equation of the line is y = -x + 2 Not in the list of answer choices which is weird.

hero (hero):

Are you sure you posted the problem correctly. Is the point (4,-2) correct, for example?